Projection Admin 10/10



To do this evening/ was completed: Matthew flipped the breaker back on for T1 (which is in 4/6 server station) for the PA speakers. I tested mics multiple times and they're definitely working and, more notably, rather hot. For intros/Q&As, we can bring down the usual level for the mics (on the cp750) from 7.0 to 6.5 or 6.0, I reckon. I did not have the time to get the massive ladder out to adjust the actual speakers, but I think it's worth doing, just so everything's a little even keel. 

I added the the Queer Media Mixer slideshow that Alec made to a folder on Plex (which looks great! I tested it!), but the T4T one (for some reason) did not export with the right size, so I'm exporting another one right now to the desktop named T4T_Mixer_133_5hrs. Not sure if it's actually 5hrs but it is certainly more than 3. In any case, it should (hopefully) be the right size. Should be done by late tonight. 

Presentation Issues:

Preshow Clips/Playlist: FYI - NF preroll loop requires 1.85 lens file and 1.78 masking. You will notice the necessary changes for that in each playlist with the preroll and or slides that also require it. Built playlists for Serpent's Skin/Lesbian Space Princess/Nīnxs/All About the T).
The new NewFest V3 trailer has been added to all shows. 

We still need to get keys for Josie and the Pussycats - this and She's the He playlists have not been built

Note: Centre spots setting for T2 has been noted on both event memos for NF this weekend. It is 502. 
I also noted the Centre spots setting for T1 shows on all NF event memos. It is 803.

Trailer Updates:

Special Events/ AV set up: Since we moved two shorts programs from T1 to T2 (this weekend's Food, Glorious Food, and In Your Face) I went through both, liberally, to make the necessary adjustments and adjusted one of the centre spots settings to fit the slide for Q&As (NF logo). I had Jess take some photos of me in a chair, etc, and put them in our thread so that NF folks are happy.  

Ordering/Receiving (posters & trailers):

Ordering/Receiving (content): Received the new trailer list from Kat today. Ingested the following: 
TLR_NWB_Strange Days 
TLR_MUBI_Lurker 
TLR_NWB_Cyber City Odeo
TLR_NPP_FORECAST_1017_A 
TLR_NPP_FORECAST_1017_B 
PS_SE_ReconsiderThis_V2_5min_OCT2025_ScoobyDoo

^please make new playlists this evening! 

Outgoing/Shipping (content): Deleted a bunch of content today. Servers are less stressed. 

Print/video Inspection:

Asset tracking doc updated/checked?:

Maintenance/cleaning:

Cue sheets updated/checked (every shift): Yes. 

Marquee updated?:

Posters updated?:

Cleaning lights on for the night?:

Projection cart by entrance?:

Additional Notes:

Comments

Popular posts from this blog

NPP Projection Open 12/28

NPP Projection Admin (Lillian/Will) 7/7

Projection Admin 9/8 Lillian/Will